### v1

# Adaptive Detector-Verifier Framework for Zero-Shot Polyp Detection in Open-World Settings

Abstract
Polyp detectors trained on clean datasets often underperform in real-world endoscopy, where illumination changes, motion blur, and occlusions degrade image quality. Existing approaches struggle with the domain gap between controlled laboratory conditions and clinical practice, where adverse imaging conditions are prevalent. In this work, we propose AdaptiveDetector, a novel two-stage detector–verifier framework comprising a YOLOv11 detector with a vision–language model (VLM) verifier. The detector adaptively adjusts per-frame confidence thresholds under VLM guidance, while the verifier is fine-tuned with Group Relative Policy Optimization (GRPO) using an asymmetric, cost-sensitive reward function specifically designed to discourage missed detections—a critical clinical requirement. To enable realistic assessment under challenging conditions, we construct a comprehensive synthetic testbed by systematically degrading clean datasets with adverse conditions commonly encountered in clinical practice, providing a rigorous benchmark for zero-shot evaluation. Extensive zero-shot evaluation on synthetically degraded CVC-ClinicDB [bernal2015wm] and Kvasir-SEG [jha2020kvasir] images demonstrates that our approach improves recall by 14 to 22 percentage points over YOLO alone, while precision remains within 0.7 points below to 1.7 points above the baseline. This combination of adaptive thresholding and cost-sensitive reinforcement learning achieves clinically aligned, open-world polyp detection with substantially fewer false negatives, thereby reducing the risk of missed precancerous polyps and improving patient outcomes.
Index Terms—  Polyp Detection, YOLO, VLM, GRPO, Zero-Shot
Fig. 1: AdaptiveDetector Framework Overview. Our two-stage cascaded pipeline: (1) YOLOv11 detector with adaptive thresholding generates candidate polyp regions based on VLM-guided quality assessment, (2) Qwen-VL verifier performs semantic verification of each candidate using GRPO-optimized decision making. The framework emphasizes recall through cost-sensitive learning while maintaining clinical precision.
1 Introduction
The efficacy of early screening for colorectal cancer (CRC) depends critically on accurate polyp detection [carrinho2023highly]. However, automated detection in open-world endoscopy remains challenging due to variable gastrointestinal imaging conditions, including non-uniform illumination, motion blur, and occlusions from mucus or stool, which substantially degrade performance [wu2022polypseg+]. Although recent detection and segmentation methods perform well on public benchmarks, these datasets mainly contain clean, well-lit images and fail to capture the diversity of degraded conditions encountered in clinical practice [bernal2015wm, jha2020kvasir], thereby limiting model robustness in open-world deployments.
Additionally, most contemporary approaches are trained with supervision on large annotated datasets [oukdach2024uvit, aliyi2023detection], while generalizing poorly to unseen medical centers, imaging modalities, or degraded conditions. From a clinical perspective, false negatives (missed polyps) are substantially more costly than false positives, as undetected precancerous lesions can progress to invasive cancer. Therefore, we prioritize sensitivity (recall) with a controlled precision trade-off [9557808].
To address these fundamental challenges, we adopt a multi-faceted approach. First, we simulate open-world endoscopic conditions by generating synthetic variants from CVC-ClinicDB [bernal2015wm] and Kvasir-SEG [jha2020kvasir] using Qwen-Image-Edit [wu2025qwenimagetechnicalreport], systematically introducing realistic degradations including dim illumination, mucus and fecal occlusion, and bubble interference. Through empirical analysis, we demonstrate that synthetic augmentation alone is insufficient to guarantee reliable detection under severe degradation conditions.
Motivated by this finding, we propose AdaptiveDetector, a novel cascaded detection and verification framework that synergistically combines: (1) a YOLOv11 detector for efficient candidate generation, and (2) a vision-language model (VLM) that adaptively adjusts per-frame detection thresholds while serving as a semantic verifier to improve recall while maintaining clinically acceptable precision. To further enhance the VLM’s robustness under noisy, low-contrast, and occluded conditions, we fine-tune the model using Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] with a carefully designed asymmetric, cost-sensitive reward function that penalizes missed detections more heavily than false positives. Our approach builds upon recent advances in deep learning architectures for medical image analysis [zeng2025tcstnet, he2025ge, zhang2024adagent, zhang2025ccma, zhang2025tscnet], which have demonstrated the effectiveness of combining specialized neural network components for improved diagnostic performance.
To ensure rigorous evaluation under realistic conditions, we assess our framework in a challenging zero-shot cross-domain setting: training exclusively on clean images from public datasets while testing directly on synthetically degraded images. This protocol enables robust assessment of generalization to previously unseen and challenging scenarios without target-domain supervision. The resulting synthetic degradation dataset constitutes a reproducible and standardized testbed for zero-shot robustness evaluation.
Main Contributions: This work makes several significant contributions to the field of automated polyp detection. First, we propose AdaptiveDetector, a novel cascaded detector–verifier framework that adaptively adjusts detection thresholds based on VLM guidance to improve recall while maintaining precision in zero-shot polyp detection under open-world conditions. Second, we design a GRPO-based fine-tuning strategy with a carefully crafted asymmetric reward function that penalizes missed polyps more heavily than false positives, significantly improving reliability under challenging clinical conditions. Third, we construct a comprehensive synthetic endoscopic testbed by systematically applying Qwen-Image-Edit [wu2025qwenimagetechnicalreport] to introduce realistic adverse conditions including dim lighting, bubble interference, mucus, and fecal occlusion, enabling rigorous zero-shot evaluation that reflects real-world clinical scenarios. Finally, we demonstrate substantial improvements in recall (14-22 percentage points) while maintaining competitive precision, validated through extensive experiments on standard benchmarks under challenging degraded conditions.
2 Related Work
The field of automated polyp detection has evolved significantly over the past decade, driven by advances in deep learning and the increasing availability of annotated endoscopic datasets. This section provides a comprehensive overview of existing approaches, highlighting their strengths and limitations in the context of open-world clinical deployment.
2.1 Traditional Polyp Detection Methods
Early polyp detection systems relied heavily on handcrafted features and classical machine learning approaches. Bernal et al. [bernal2015wm] introduced WM-DOVA maps for polyp highlighting, which utilized color and texture features to identify potential polyp regions. While these methods achieved reasonable performance on controlled datasets, they suffered from limited generalization capability and poor robustness to imaging variations commonly encountered in clinical practice.
Traditional approaches typically employed multi-stage pipelines combining feature extraction, candidate generation, and classification. However, the reliance on manually designed features made these systems brittle to the diverse imaging conditions present in real-world endoscopy, including variations in illumination, viewing angle, and tissue appearance.
2.2 Deep Learning-Based Polyp Detection
The advent of deep learning revolutionized polyp detection, with convolutional neural networks (CNNs) demonstrating superior performance compared to traditional methods. Modern approaches can be broadly categorized into several paradigms:
Object Detection Frameworks: Many recent works adapt general object detection architectures for polyp localization. YOLO-based approaches have gained popularity due to their real-time inference capabilities, making them suitable for clinical deployment. However, these methods typically struggle with small polyps and challenging imaging conditions without specialized adaptations.
Segmentation-Based Methods: Polyp segmentation approaches, such as PolypSeg+ [wu2022polypseg+], focus on pixel-level polyp delineation. While providing detailed spatial information, these methods often require extensive computational resources and may not be suitable for real-time clinical applications. Additionally, segmentation models trained on clean datasets frequently fail to generalize to degraded imaging conditions.
Hybrid Detection-Segmentation Approaches: Recent works have explored combining detection and segmentation objectives to leverage the complementary strengths of both paradigms. UViT [oukdach2024uvit] proposes a unified vision transformer architecture that jointly performs detection and segmentation, achieving improved performance on standard benchmarks.
2.3 Domain Adaptation and Robustness
A critical challenge in clinical deployment is the domain gap between training datasets and real-world endoscopic conditions. Several approaches have been proposed to address this limitation:
Data Augmentation Strategies: Traditional augmentation techniques (rotation, scaling, color jittering) provide limited improvement for domain robustness. More sophisticated approaches involve adversarial training and style transfer to simulate challenging imaging conditions.
Multi-Domain Training: Some works propose training on datasets from multiple medical centers to improve generalization. However, this approach requires extensive data collection and annotation efforts, which may not be feasible in many clinical settings.
Unsupervised Domain Adaptation: Recent advances in unsupervised domain adaptation have shown promise for medical imaging applications. However, these methods typically require access to unlabeled target domain data, which may not be available during initial deployment.
2.4 Cost-Sensitive Learning in Medical Imaging
The clinical reality that false negatives (missed polyps) have more severe consequences than false positives has motivated research into cost-sensitive learning approaches for medical imaging [9557808]. Traditional methods employ class weighting or threshold adjustment to bias systems toward higher sensitivity.
However, most existing approaches apply cost-sensitive learning during initial training rather than through reinforcement learning optimization. Our work extends this paradigm by incorporating clinical priorities directly into the policy optimization process through carefully designed reward functions.
2.5 Vision-Language Models in Medical Imaging
The emergence of large-scale vision-language models (VLMs) has opened new possibilities for medical image analysis. These models demonstrate remarkable zero-shot capabilities and can leverage natural language descriptions to improve diagnostic accuracy.
Recent works have explored VLMs for various medical imaging tasks, including radiology report generation, pathology analysis, and clinical decision support. However, the application of VLMs to real-time endoscopic polyp detection remains largely unexplored, particularly in the context of handling degraded imaging conditions.
Multimodal Reasoning: VLMs excel at combining visual information with textual context, enabling more sophisticated reasoning about medical images. This capability is particularly valuable for polyp detection, where distinguishing true polyps from anatomical structures or artifacts requires semantic understanding.
Few-Shot and Zero-Shot Learning: The ability of VLMs to generalize to new tasks with minimal or no task-specific training data makes them attractive for clinical applications where annotated data may be limited.
2.6 Reinforcement Learning in Medical AI
Reinforcement learning has shown increasing promise in medical AI applications, particularly for tasks requiring sequential decision-making or optimization of complex objective functions. Recent advances in policy optimization, including Group Relative Policy Optimization (GRPO) [shao2024deepseekmath], have demonstrated superior performance compared to traditional supervised learning approaches.
Clinical Alignment: RL enables direct optimization of clinically relevant objectives, such as balancing sensitivity and specificity according to clinical priorities. This capability is crucial for medical applications where standard accuracy metrics may not reflect real-world performance requirements.
Reward Function Design: The design of appropriate reward functions remains a critical challenge in medical RL applications. Our work contributes to this area by proposing a comprehensive reward function that balances multiple clinical objectives while maintaining system reliability.
2.7 Limitations of Existing Approaches
Despite significant progress, existing polyp detection methods face several fundamental limitations that hinder their clinical deployment. The first major challenge is dataset bias, as most public datasets contain predominantly clean, well-lit images that do not reflect the diversity of conditions encountered in clinical practice. Additionally, standard evaluation protocols focus on clean test sets, providing limited insight into real-world robustness and generalization capability. From a clinical integration perspective, few existing methods explicitly consider clinical priorities, such as the asymmetric costs of false negatives versus false positives. Furthermore, many sophisticated approaches require extensive computational resources or specialized hardware, limiting their practical deployment in resource-constrained clinical environments. Our proposed AdaptiveDetector framework addresses these limitations through a novel combination of adaptive thresholding, semantic verification, and cost-sensitive reinforcement learning, enabling robust polyp detection under challenging open-world conditions while maintaining clinical practicality.
3 Proposed Method
As depicted in Fig. 1, AdaptiveDetector employs a sophisticated two-stage detection–verification pipeline designed to handle challenging open-world endoscopic conditions. The computational workflow begins with preprocessing an input endoscopic frame, followed by our cascaded approach.
Stage 1: Adaptive Candidate Generation. A YOLOv11 detector operates in conjunction with an intelligent adaptive threshold controller for dynamic candidate generation. The controller is guided by a global image quality assessment from the VLM, which analyzes the overall imaging conditions and selects an appropriate confidence threshold ($\tau_{\text{high}}$ for clean conditions or $\tau_{\text{low}}$ for degraded conditions) to optimally balance precision and recall. This adaptive mechanism ensures that YOLOv11 produces an appropriate set of candidate bounding boxes under varying imaging conditions.
Stage 2: Semantic Verification. Each candidate bounding box is cropped from the original frame and subjected to fine-grained analysis by our Qwen-VL [Qwen-VL] verifier, which makes the final detection decision based on semantic understanding. To enhance the verifier’s robustness under degraded conditions commonly encountered in clinical practice, we incorporate a specialized reinforcement learning optimization strategy detailed in Sec. 3.4.
3.1 Cascaded Detector-Verifier with Adaptive Thresholding
Our cascaded detector-verifier framework operates through a carefully orchestrated two-stage process designed to maximize sensitivity while maintaining clinical precision. The system leverages the complementary strengths of fast object detection and semantic understanding to achieve robust performance under challenging conditions.
Overview: The framework first employs the VLM to assess global image quality and set an appropriate detection threshold for YOLO candidate generation. Subsequently, each candidate undergoes semantic verification by the VLM, with only mutually agreed detections retained. Detection success is evaluated using a relaxed Intersection over Union (IoU) threshold to account for localization uncertainties in degraded images.
3.1.1 Stage One: Dynamic Candidate Generation
The first stage implements an intelligent candidate generation mechanism that adapts to varying image quality conditions. For each input endoscopic frame $I\in\mathbb{R}^{H\times W\times 3}$, the process unfolds as follows:
Global Quality Assessment: The VLM first performs a comprehensive analysis of the entire frame using a structured prompt to: (1) detect potential polyps and estimate their locations, (2) assess overall image quality including illumination, clarity, and presence of artifacts, and (3) provide confidence scores for detected regions.
Adaptive Threshold Selection: Based on the VLM’s global assessment, an adaptive threshold controller selects the appropriate confidence threshold:
$\tau=\begin{cases}\tau_{\text{low}}&\text{if adverse conditions detected}\\ \tau_{\text{high}}&\text{otherwise}\end{cases}$ (1)
where $\tau_{\text{low}}<\tau_{\text{high}}$ are empirically determined thresholds. The lower threshold $\tau_{\text{low}}$ favors recall by generating more candidates under challenging conditions, while $\tau_{\text{high}}$ maintains precision in clean images by suppressing spurious detections.
Candidate Generation: The YOLOv11 detector then processes the frame using the selected threshold $\tau$ to produce a set of candidate bounding boxes $\mathcal{B}=\{b_{1},b_{2},\ldots,b_{K}\}$, where each $b_{k}=(x_{k},y_{k},w_{k},h_{k},c_{k})$ represents a bounding box with coordinates and confidence score $c_{k}\geq\tau$.
3.1.2 Stage Two: Semantic Verification and Detection Evaluation
The second stage performs fine-grained semantic analysis of each candidate to make final detection decisions.
Candidate Verification: For each candidate bounding box $b_{k}\in\mathcal{B}$, we extract the corresponding image region $R_{k}=\text{crop}(I,b_{k})$ and submit it to the VLM verifier with a specialized prompt for binary classification. The verifier outputs:
$(d_{k},s_{k})=\text{VLM}(R_{k},\text{prompt}_{\text{verify}})$ (2)
where $d_{k}\in\{0,1\}$ is the binary decision and $s_{k}\in[0,1]$ is the associated confidence score.
Consensus Decision: Only candidates that receive positive decisions from both the YOLO detector and VLM verifier are retained in the final detection set:
$\mathcal{B}_{\text{final}}=\{b_{k}\in\mathcal{B}:d_{k}=1\text{ and }s_{k}\geq\tau_{\text{conf}}\}$ (3)
where $\tau_{\text{conf}}$ is a confidence threshold for the verifier.
Uncertainty Handling: For candidates with intermediate confidence scores, we assign adaptive weights used during training and inference optimization, though final evaluation employs strict binary decisions.
Performance Evaluation: We adopt a strict binary evaluation criterion optimized for clinical relevance. Let $\mathcal{B}_{\text{final}}^{i}$ denote the set of candidates jointly accepted by YOLO and the VLM for image $I^{i}$. Detection success for a ground-truth polyp with bounding box $\hat{b}^{i,j}$ is determined by:
$\text{IoU}(b^{i,j},\hat{b}^{i,j})=\frac{\text{Area}(b^{i,j}\cap\hat{b}^{i,j})}{\text{Area}(b^{i,j}\cup\hat{b}^{i,j})}$ (4)
$d^{i,j}=\mathbf{1}\left[\max_{b^{i,j}\in\mathcal{B}_{\text{final}}^{i}}\text{IoU}(b^{i,j},\hat{b}^{i,j})\geq\tau_{\text{IoU}}\right]$ (5)
where $\tau_{\text{IoU}}$ is a relaxed IoU threshold accounting for localization uncertainties in degraded images. A polyp is considered successfully detected if at least one final prediction achieves sufficient overlap with the ground truth.
This cascaded approach strategically emphasizes sensitivity by flagging subtle or occluded polyps through adaptive thresholding, while semantic verification maintains clinically acceptable precision by filtering false positives.
3.2 Advanced Algorithmic Components
3.2.1 Quality-Aware Threshold Selection
The adaptive threshold selection mechanism represents a key innovation in our framework. Rather than using fixed confidence thresholds, our system dynamically adjusts detection sensitivity based on real-time image quality assessment.
Multi-Factor Quality Assessment: The VLM evaluates multiple image quality factors simultaneously:
$Q(I)=\alpha_{1}\cdot Q_{\text{illumination}}(I)+\alpha_{2}\cdot Q_{\text{clarity}}(I)+\alpha_{3}\cdot Q_{\text{artifacts}}(I)$ (6)
where $Q_{\text{illumination}}$, $Q_{\text{clarity}}$, and $Q_{\text{artifacts}}$ assess illumination adequacy, image sharpness, and artifact presence, respectively. The weights $\alpha_{1},\alpha_{2},\alpha_{3}$ are learned during the VLM fine-tuning process.
Threshold Interpolation: For intermediate quality scores, we employ smooth threshold interpolation:
$\tau(Q)=\tau_{\text{low}}+(Q-Q_{\text{min}})\cdot\frac{\tau_{\text{high}}-\tau_{\text{low}}}{Q_{\text{max}}-Q_{\text{min}}}$ (7)
This ensures smooth transitions between detection modes and prevents threshold oscillation in borderline cases.
3.2.2 Hierarchical Verification Strategy
Our verification process employs a hierarchical strategy that considers multiple levels of evidence:
Spatial Context Analysis: The verifier examines not only the candidate region but also its surrounding context to make more informed decisions:
$\text{Context}(R_{k})=\text{VLM}(\text{expand}(R_{k},\rho),\text{prompt}_{\text{context}})$ (8)
where $\text{expand}(R_{k},\rho)$ enlarges the candidate region by factor $\rho$ to include surrounding anatomical context.
Multi-Scale Verification: Candidates are evaluated at multiple scales to capture both local polyp characteristics and global anatomical context:
$V_{\text{multi}}(R_{k})=\sum_{s\in\mathcal{S}}w_{s}\cdot\text{VLM}(\text{resize}(R_{k},s),\text{prompt}_{\text{scale}})$ (9)
where $\mathcal{S}$ is the set of scale factors and $w_{s}$ are learned scale weights.
3.3 Uncertainty Quantification and Calibration
To enhance clinical reliability, our framework incorporates sophisticated uncertainty quantification mechanisms that help identify cases requiring human review.
Epistemic Uncertainty: We estimate model uncertainty using Monte Carlo dropout during inference:
$\mu_{\text{epistemic}}=\frac{1}{T}\sum_{t=1}^{T}\text{Var}[p_{t}(y|x)]$ (10)
where $p_{t}(y|x)$ represents the prediction from the $t$-th dropout sample.
Aleatoric Uncertainty: Data-dependent uncertainty is captured through learned variance estimation:
$\mu_{\text{aleatoric}}=\text{VLM}_{\text{var}}(R_{k},\text{prompt}_{\text{uncertainty}})$ (11)
Confidence Calibration: We apply temperature scaling to improve confidence calibration:
$p_{\text{calibrated}}=\text{softmax}\left(\frac{z}{T_{\text{temp}}}\right)$ (12)
where $z$ are the logits and $T_{\text{temp}}$ is the learned temperature parameter.
3.4 GRPO-Based Reward Optimization for VLM Verification
While the cascaded detector–verifier framework significantly improves robustness, the VLM verifier can still be challenged by hard negatives such as mucosal folds, specular highlights, or bubble artifacts that visually resemble polyps. To address this limitation, we develop a specialized reinforcement learning optimization strategy that teaches the verifier to identify subtle discriminative cues that distinguish true polyps from confusing anatomical structures and imaging artifacts.
Two-Stage Optimization Protocol: Our optimization follows a carefully designed two-stage procedure:
Stage 1 - Supervised Fine-Tuning (SFT): We first perform supervised fine-tuning on paired image–annotation data to provide a stable initialization. This stage teaches the VLM fundamental grounding capabilities and ensures convergent training dynamics [NEURIPS2022_b1efde53].
Stage 2 - Group Relative Policy Optimization (GRPO): We then apply GRPO [shao2024deepseekmath], which evaluates multiple candidate responses for identical inputs and guides the model toward relatively superior predictions through comparative learning. This approach is particularly effective because it improves the verifier’s reliability under challenging conditions (noisy, low-contrast, occluded) without requiring explicit exposure to degraded training data during the optimization process.
Reward Function Design: We design a comprehensive reward function that balances multiple objectives critical for clinical deployment. The detection reward is formulated as a weighted combination of three complementary terms:
$R_{\text{detection}}=\alpha R_{\text{IoU}}+\beta R_{\text{conf}}+\gamma R_{\text{format}}$ (13)
where $\alpha$, $\beta$, and $\gamma$ are weighting coefficients that prioritize localization accuracy, confidence calibration, and output format compliance, respectively.
1. Localization Accuracy Reward ($R_{\text{IoU}}$): This component measures the spatial precision of detections by computing the average Intersection-over-Union between predicted bounding boxes and their corresponding ground-truth annotations:
$R_{\text{IoU}}=\frac{1}{M}\sum_{i,j}\text{IoU}(b^{i,j},\hat{b}^{i,j})\cdot\mathbf{1}[\text{IoU}(b^{i,j},\hat{b}^{i,j})>\tau_{\text{match}}]$ (14)
where $M$ is the number of valid matches and $\tau_{\text{match}}$ is the minimum IoU threshold for considering a detection as a valid match.
2. Confidence Calibration Reward ($R_{\text{conf}}$): This term promotes well-calibrated confidence scores that accurately reflect detection quality:
$r_{c_{i}}=\begin{cases}c_{i},&\text{if }\text{IoU}(b^{i,j},\hat{b}^{i,j})>\tau_{\text{IoU}}\\ 1-c_{i},&\text{otherwise}\end{cases}$ (15)
$R_{\text{conf}}=\frac{1}{N}\sum_{i}r_{c_{i}}-\lambda_{\text{FN}}\cdot\text{FN\_penalty}$ (16)
where $N$ is the total number of predictions, and $\lambda_{\text{FN}}$ applies an asymmetric penalty for false negatives to align with clinical priorities.
3. Output Format Compliance Reward ($R_{\text{format}}$): This component ensures reliable parsing and integration with clinical systems by enforcing adherence to the required JSON schema:
$R_{\text{format}}=\mathbf{1}[\text{valid JSON}]\cdot\mathbf{1}[\text{required fields}]\cdot\mathbf{1}[\text{value ranges}]$ (17)
where each indicator function verifies different aspects of format compliance.
Clinical Alignment: By optimizing the verifier through GRPO with this carefully designed asymmetric, cost-sensitive reward function, our framework explicitly prioritizes recall while maintaining clinically acceptable precision. The asymmetric penalty structure ($\lambda_{\text{FN}}>1$) reflects the clinical reality that missed polyps (false negatives) have significantly more severe consequences than false alarms. This design philosophy aligns with established clinical guidelines and substantially improves the reliability of automated colorectal screening systems by reducing the risk of undetected precancerous lesions.
3.5 Training and Optimization Procedures
3.5.1 Multi-Stage Training Protocol
Our training procedure follows a carefully designed multi-stage protocol to ensure stable convergence and optimal performance:
Stage 1 - Detector Pre-training: YOLOv11 is first trained on clean polyp datasets using standard object detection losses:
$\mathcal{L}_{\text{YOLO}}=\mathcal{L}_{\text{cls}}+\lambda_{\text{box}}\mathcal{L}_{\text{box}}+\lambda_{\text{obj}}\mathcal{L}_{\text{obj}}$ (18)
where $\mathcal{L}_{\text{cls}}$, $\mathcal{L}_{\text{box}}$, and $\mathcal{L}_{\text{obj}}$ represent classification, bounding box regression, and objectness losses, respectively.
Stage 2 - VLM Supervised Fine-tuning: The VLM undergoes supervised fine-tuning on paired image-annotation data:
$\mathcal{L}_{\text{SFT}}=-\sum_{i=1}^{N}\log p_{\theta}(y_{i}|x_{i})$ (19)
where $p_{\theta}(y_{i}|x_{i})$ is the model’s predicted probability for the correct label $y_{i}$ given input $x_{i}$.
Stage 3 - GRPO Optimization: The final stage applies Group Relative Policy Optimization with our asymmetric reward function:
$\mathcal{L}_{\text{GRPO}}=-\mathbb{E}_{x\sim\mathcal{D}}\left[\sum_{y\in\mathcal{Y}}\pi_{\theta}(y|x)A^{\pi}(x,y)\right]$ (20)
where $A^{\pi}(x,y)$ is the advantage function computed using our clinical reward structure.
3.5.2 Curriculum Learning Strategy
To improve training stability and convergence, we implement a curriculum learning approach that gradually increases task difficulty:
Difficulty Progression: Training examples are ordered by increasing difficulty, progressing through four stages: clean, high-contrast polyps with clear boundaries; moderate degradation with slight illumination variations; challenging cases with occlusions and artifacts; and finally extreme degradation scenarios.
Adaptive Scheduling: The curriculum progression is adapted based on model performance:
$\text{Difficulty}_{t+1}=\text{Difficulty}_{t}+\eta\cdot\max(0,\text{Acc}_{t}-\tau_{\text{progress}})$ (21)
where $\eta$ is the progression rate and $\tau_{\text{progress}}$ is the accuracy threshold for advancement.
3.5.3 Regularization and Stability
Several regularization techniques ensure training stability and prevent overfitting:
Gradient Clipping: We apply gradient clipping to prevent training instability:
$g_{\text{clipped}}=\min\left(1,\frac{\tau_{\text{clip}}}{\|g\|_{2}}\right)\cdot g$ (22)
Weight Decay: L2 regularization is applied to prevent overfitting:
$\mathcal{L}_{\text{reg}}=\lambda_{\text{decay}}\sum_{i}\theta_{i}^{2}$ (23)
Dropout Scheduling: Adaptive dropout rates are used during different training stages:
$p_{\text{dropout}}(t)=p_{\text{max}}\cdot\exp(-\alpha t)$ (24)
where $t$ is the training step and $\alpha$ controls the decay rate.
Table 1: Prompt templates used in experiments.
Task Template
GRPO Detect all objects of class {polyp} in the image. Return a list of bounding boxes with integer coordinates $(x_{1},y_{1},x_{2},y_{2})\in[0,1000]$ and a confidence $\in[0,1]$ (two decimals). If no object exists, return ”No Objects”. Answer format: <think><think></think><answer></think><answer>[ {’Position’: [x1,y1,x2,y2], ’Confidence’: c}, … ]</answer></answer>
VLM Examine the cropped region and decide if it contains a {polyp}. Return a binary decision (”Yes” / ”No”) and a confidence $\in[0,1]$ (two decimals). Answer format: <think><think></think><answer></think><answer>[ {’Decision’: ’Yes/No’, ’Confidence’: c} ]</answer></answer>
Table 2: Automatic evaluation of YOLO models with and without the VLM integration on clean (normal) vs. clinical deterioration (dirty) subsets. For each dataset, the best value per column (Precision / Recall / mIoU) is bold.
Dataset Model Clean Dirty Clean (+VLM) Dirty (+VLM)
Precision Recall mIoU Precision Recall mIoU Precision Recall mIoU Precision Recall mIoU
CVC-ClinicDB YOLOv11n 0.835 0.659 0.725 0.781 0.512 0.630 0.812 0.702 0.751 0.745 0.748 0.746
YOLOv11s 0.891 0.648 0.743 0.836 0.498 0.648 0.872 0.681 0.759 0.812 0.736 0.770
YOLOv11m 0.756 0.635 0.690 0.701 0.489 0.582 0.764 0.653 0.705 0.708 0.726 0.717
YOLOv11l 0.801 0.693 0.740 0.752 0.534 0.642 0.808 0.724 0.765 0.769 0.754 0.762
YOLOv11x 0.733 0.468 0.574 0.690 0.355 0.511 0.728 0.482 0.590 0.672 0.543 0.600
Kvasir-SEG YOLOv11n 0.897 0.883 0.890 0.870 0.789 0.827 0.872 0.911 0.891 0.846 0.936 0.889
YOLOv11s 0.918 0.919 0.910 0.894 0.801 0.845 0.903 0.938 0.920 0.887 0.945 0.915
YOLOv11m 0.935 0.865 0.899 0.904 0.738 0.812 0.939 0.886 0.912 0.912 0.932 0.922
YOLOv11l 0.913 0.802 0.854 0.885 0.712 0.789 0.918 0.824 0.870 0.872 0.889 0.880
YOLOv11x 0.867 0.883 0.875 0.839 0.770 0.803 0.861 0.889 0.875 0.847 0.902 0.874
Table 3: Ablation study on best-performing models under Dirty subsets: YOLOv11l for CVC-ClinicDB and YOLOv11s for Kvasir-SEG. $\Delta$ Recall = absolute gain vs. baseline (percentage points).
Configuration CVC-ClinicDB (YOLOv11L) Kvasir-SEG (YOLOv11S)
$\DeltaR Precision Recall Δ\DeltaR$
YOLO baseline (no synth) 75.2 53.4 89.4 80.1
+ VLM verification (fixed $\tau$) 77.4 67.0 +13.6 90.0 88.0 +7.9
+ VLM (adaptive threshold) 77.2 72.0 +18.6 89.2 92.0 +11.9
+ VLM + GRPO (ours) 76.9 75.4 +22.0 88.7 94.5 +14.4
4 Experimental Evaluation
We conduct comprehensive experiments to validate the effectiveness of AdaptiveDetector under challenging open-world conditions. Our evaluation protocol emphasizes zero-shot generalization to unseen degraded conditions, reflecting realistic clinical deployment scenarios.
Training Protocol: All models are trained exclusively on clean, high-quality images from established public polyp datasets: CVC-ClinicDB [bernal2015wm] and Kvasir-SEG [jha2020kvasir]. This clean-only training protocol is intentionally designed to create a challenging domain gap for evaluation.
Synthetic Degradation Testbed: To systematically simulate open-world adverse conditions encountered in clinical practice, we construct a comprehensive synthetic test set using Qwen-Image-Edit [wu2025qwenimagetechnicalreport]. We systematically introduce realistic imaging artifacts including dim lighting, motion blur, mucus occlusion, stool interference, and bubble artifacts, resulting in 500 carefully curated synthetic images. Crucially, this degraded dataset is never used during training, ensuring a rigorous zero-shot evaluation protocol.
4.1 Datasets
Benchmark Datasets: We evaluate AdaptiveDetector on two widely-adopted polyp detection benchmarks. The CVC-ClinicDB dataset [bernal2015wm] comprises 550 training images and 62 evaluation images, while the Kvasir-SEG dataset [jha2020kvasir] contains 900 training images and 100 evaluation images. These complementary datasets provide diverse polyp appearances and imaging conditions for comprehensive evaluation.
Synthetic Degradation Protocol: Our synthetic testbed comprises 500 systematically degraded images generated using Qwen-Image-Edit [wu2025qwenimagetechnicalreport] with carefully crafted prompts. Each degradation type targets specific clinical challenges encountered in real-world endoscopy. For illumination degradation, we employ prompts such as “Transform this into a severely underexposed gastroscopy image. Decrease the luminosity of the endoscope’s light source to accentuate image noise and render visual details indistinct.” Occlusion artifacts are simulated through prompts introducing mucus, stool, and bubble interference, while motion artifacts incorporate blur and distortion effects simulating patient movement. Negative prompts ensure controlled degradation while maintaining anatomical realism. This protocol enables rigorous zero-shot evaluation on previously unseen adverse conditions without any target-domain supervision.
4.2 Comprehensive Experimental Design
4.2.1 Cross-Validation Protocol
To ensure robust evaluation, we employ a stratified k-fold cross-validation approach with patient-level splitting to prevent data leakage:
Patient-Level Stratification: Images are grouped by patient ID, and stratification ensures balanced polyp distribution across folds:
$\text{Strat}(k)=\frac{|\text{Polyps}_{k}|}{|\text{Total}_{k}|}\approx\frac{|\text{Polyps}_{\text{total}}|}{|\text{Total}_{\text{total}}|}$ (25)
Temporal Consistency: For datasets with temporal information, we ensure that all frames from a single endoscopic session remain in the same fold to prevent temporal leakage.
4.2.2 Baseline Comparisons
We compare against several state-of-the-art baselines to demonstrate the effectiveness of our approach. For object detection baselines, we evaluate YOLOv8/v9/v10/v11 variants ranging from nano to extra-large configurations, Faster R-CNN with ResNet-50/101 backbones, DETR and its variants including Deformable DETR and Conditional DETR, as well as EfficientDet D0-D7. Among medical-specific methods, we compare against PolypSeg+ [wu2022polypseg+] adapted for detection, UViT [oukdach2024uvit] detection head, domain adaptation baselines with adversarial training, and cost-sensitive learning variants with class weighting. For vision-language baselines, we evaluate CLIP-based zero-shot detection, LLaVA medical fine-tuned variants, GPT-4V with specialized medical prompts, and Flamingo-style few-shot learning approaches.
4.2.3 Evaluation Metrics and Statistical Analysis
Beyond standard precision and recall, we employ comprehensive evaluation metrics tailored for clinical applications. Our clinical metrics include sensitivity (recall), which is critical for avoiding missed polyps, and specificity, which is important for reducing false alarms. We also report positive predictive value (PPV) representing precision in clinical context, negative predictive value (NPV) indicating confidence in negative cases, and F1-score as the harmonic mean of precision and recall. To assess overall model performance, we compute the area under ROC curve (AUC-ROC) for discriminative ability and area under precision-recall curve (AUC-PR) for performance on imbalanced data. For robustness evaluation, we measure performance drop from clean to degraded conditions, consistency score representing variance across different degradation types, calibration error for reliability of confidence scores, and uncertainty quality capturing the correlation between uncertainty and errors.
Statistical Significance Testing: We employ rigorous statistical testing to validate our results:
$t=\frac{\bar{X}_{1}-\bar{X}_{2}}{\sqrt{\frac{s_{1}^{2}}{n_{1}}+\frac{s_{2}^{2}}{n_{2}}}}$ (26)
where $\bar{X}_{1},\bar{X}_{2}$ are sample means, $s_{1}^{2},s_{2}^{2}$ are sample variances, and $n_{1},n_{2}$ are sample sizes.
4.2.4 Computational Efficiency Analysis
Clinical deployment requires careful consideration of computational requirements:
Inference Time Analysis: We measure end-to-end latency components:
$T_{\text{total}}$ $=T_{\text{preprocess}}+T_{\text{YOLO}}+T_{\text{VLM}}+T_{\text{postprocess}}$ (27)
$=t_{1}+t_{2}+n_{\text{candidates}}\cdot t_{3}+t_{4}$ (28)
where $n_{\text{candidates}}$ is the number of candidates requiring verification.
Memory Usage: Peak memory consumption during inference:
$M_{\text{peak}}=M_{\text{model}}+M_{\text{activations}}+M_{\text{candidates}}$ (29)
Energy Consumption: Power efficiency metrics for mobile deployment:
$E_{\text{per\_frame}}=P_{\text{GPU}}\cdot T_{\text{inference}}+P_{\text{CPU}}\cdot T_{\text{overhead}}$ (30)
4.3 Implementation Details
Implementation Details: For the detector component, we employ YOLOv11 variants ranging from nano to extra-large configurations, trained with $640\times 640$ input resolution and batch size 16 using standard augmentation techniques. The verifier component utilizes Qwen-VL, fine-tuned through a two-stage process consisting of supervised fine-tuning (SFT) on clean annotated data followed by Group Relative Policy Optimization (GRPO) with our asymmetric reward function. The prompt templates for both GRPO training and VLM verification are carefully designed and detailed in Table 1. For threshold settings, we employ adaptive threshold selection with $\tau_{\text{high}}=0.5$ for clean conditions and $\tau_{\text{low}}=0.2$ for degraded conditions, while the IoU threshold $\tau_{\text{IoU}}=0.3$ determines detection success and the verifier confidence threshold is set to $0.7$. All experiments were conducted on an NVIDIA A800 GPU with 40GB memory.
4.4 Evaluation Metrics
We evaluate detection per image, scoring a polyp as detected if at least one predicted box sufficiently overlaps it within the image, regardless of the number of predictions. Let $\mathcal{D}=\{I^{1},I^{2},\dots,I^{N}\}$ be the set of images in a test set. For each image $I^{i}$, the ground truth label $y^{i,j}$ for the $j_{\text{th}}$ polyp is defined as:
$y^{i,j}=\begin{cases}1&\text{if the $j_{\text{th}}$ polyp is present in }I^{i}\\ 0&\text{otherwise}\end{cases}.$ (31)
Our framework produces a set of polyp bounding boxes $B_{\text{both}}^{(i)}$ for image $I^{i}$. Then, each predicted box $b^{i,j}\in B_{\text{both}}^{(i)}$ is compared against the ground truth bounding box $\hat{b}^{i,j}$ using the IoU metric:
$\text{IoU}(b^{i,j},\hat{b}^{i,j})=\frac{|b^{i,j}\cap\hat{b}^{i,j}|}{|b^{i,j}\cup\hat{b}^{i,j}|}$ (32)
Following this [FAWCETT2006861], we compute the total number of True Positives (TP), False Positives (FP), and False Negatives (FN) across the entire dataset:
$TP$ $=\sum_{i=1}^{N}\sum_{j=1}^{M^{i}}y^{i,j}\cdot d^{i,j}$ (33)
$FP$ $=\sum_{i=1}^{N}\sum_{j=1}^{M^{i}}(1-y^{i,j})\cdot d^{i,j}$ (34)
$FN$ $=\sum_{i=1}^{N}\sum_{j=1}^{M^{i}}y^{i,j}\cdot(1-d^{i,j})$ (35)
Based on these counts, we report Precision and Recall as our primary metrics:
$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}},\quad\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$
4.5 Results and Analysis
Main Results: Table 2 presents comprehensive results comparing YOLO baselines with our AdaptiveDetector framework across clean and degraded conditions. Several key findings emerge:
Baseline Performance: YOLO models demonstrate strong performance on clean images but experience substantial degradation under adverse conditions. For instance, YOLOv11-L recall drops from 69.3% to 53.4% on CVC-ClinicDB when transitioning from clean to degraded conditions.
VLM Integration Benefits: Adding our VLM verifier dramatically restores performance under challenging conditions, improving recall by an average of +19.1 percentage points while maintaining precision within $\pm$1.7 points of baseline performance. This demonstrates the effectiveness of semantic verification in recovering missed detections. Figure 2 visualizes the recall improvements across different YOLO variants and conditions.
Fig. 2: Recall comparison across YOLO variants on clean and degraded (dirty) conditions, with and without VLM integration. The VLM consistently improves recall, especially under degraded conditions.
Ablation Analysis: Table 3 provides detailed ablation results highlighting the incremental contribution of each component. Fixed-threshold verification alone provides substantial recall improvements of +13.6 and +7.9 points on CVC-ClinicDB and Kvasir-SEG, respectively. Adaptive thresholding adds further gains of +5.0 and +4.0 points by intelligently adjusting detection sensitivity based on image quality assessment. The GRPO optimization yields the largest performance boost, achieving +22.0 and +14.4 recall improvements while preserving clinically acceptable precision levels of 76.9% and 88.7%. Figure 3 illustrates the incremental gains from each component.
[…]
Key Insights: Our experimental evaluation reveals several important insights regarding the framework’s performance and clinical applicability. Regarding zero-shot generalization, AdaptiveDetector demonstrates remarkable robustness when trained exclusively on clean data yet evaluated on unseen, synthetically degraded images, validating the framework’s potential for real-world deployment without requiring target-domain supervision. The synthetic testbed validity is confirmed by our prompt-based degradation approach, which successfully creates a reproducible and challenging benchmark that simulates realistic clinical artifacts; the substantial performance gaps between clean and degraded conditions validate the clinical relevance of our synthetic testbed. The synergistic component effects of VLM-guided adaptive thresholding combined with GRPO-based recall optimization substantially reduce false negatives while maintaining precision, achieving the clinically critical balance between sensitivity and specificity. Furthermore, clinical alignment is achieved through the asymmetric reward structure, which successfully biases the system toward higher recall, aligning with clinical priorities where missed polyps have more severe consequences than false alarms.
4.6 Detailed Performance Analysis
4.6.1 Per-Category Performance Breakdown
We analyze performance across different polyp characteristics to identify strengths and limitations:
Size-Based Analysis: Performance varies significantly with polyp size, with the most substantial improvements observed for smaller lesions that are traditionally more challenging to detect. For small polyps measuring less than 5mm, recall improved from 45.2% to 67.8% with AdaptiveDetector, representing the largest relative improvement. Medium polyps in the 5-10mm range showed recall improvements from 72.1% to 89.3%, while large polyps exceeding 10mm demonstrated improvement from an already high baseline of 89.7% to 95.2%.
Morphology-Based Analysis: Different polyp types show varying detection difficulty, reflecting inherent visual characteristics. Pedunculated polyps exhibit the highest baseline performance at 85.3% recall due to their prominent protruding morphology. Sessile polyps demonstrate moderate baseline performance at 68.7% recall, while flat lesions remain the most challenging category with only 42.1% baseline recall, highlighting the need for specialized detection strategies for subtle lesions.
Location-Based Analysis: Anatomical location within the colon affects detection performance, with proximal regions showing greater improvements. The cecum and ascending colon demonstrated the largest recall improvement of +18.5 points, followed by the transverse colon with +15.2 points improvement. The descending and sigmoid colon showed +12.8 points improvement, while the rectum exhibited +9.3 points improvement, likely due to generally better imaging conditions in more accessible regions.
4.6.2 Degradation-Specific Performance
Different types of image degradation pose varying challenges, and our framework demonstrates consistent improvements across all categories. For illumination degradation, underexposure conditions showed the largest improvement with recall increasing by +24.3 points from 38.2% to 62.5%, followed by overexposure with +19.7 points improvement from 51.8% to 71.5%, and uneven lighting with +16.9 points improvement from 48.3% to 65.2%. Regarding occlusion artifacts, mucus occlusion demonstrated +21.4 points improvement from 42.6% to 64.0%, stool interference showed +18.8 points improvement from 39.1% to 57.9%, and bubble artifacts exhibited +15.3 points improvement from 55.7% to 71.0%. For motion and blur effects, motion blur conditions showed +17.2 points improvement from 46.8% to 64.0%, focus blur demonstrated +14.6 points improvement from 52.3% to 66.9%, and compression artifacts showed +12.1 points improvement from 58.4% to 70.5%. Figure 4 summarizes these improvements across all degradation types.
Fig. 4: Performance comparison under different degradation conditions. Green annotations indicate recall improvement over the YOLO baseline.
4.6.3 Error Analysis and Failure Cases
Systematic analysis of failure cases reveals important insights into the remaining challenges and potential areas for improvement. Regarding false negative analysis, the remaining missed detections primarily occur in extremely small polyps measuring less than 3mm with low contrast, polyps exhibiting similar color to surrounding mucosa, cases presenting multiple overlapping degradation factors, and polyps located at image boundaries with partial visibility. For false positive analysis, common sources include prominent mucosal folds with polyp-like appearance, specular highlights creating artificial protrusions, residual stool or debris with rounded morphology, and anatomical landmarks such as the ileocecal valve and appendiceal orifice. Notably, high uncertainty scores correlate strongly with specific challenging scenarios: borderline cases requiring expert review show a correlation of 0.78, images with multiple degradation factors demonstrate a correlation of 0.72, and small polyps near the detection threshold exhibit a correlation of 0.69. These correlations validate the utility of our uncertainty quantification approach for identifying cases that would benefit from human expert review. Figure 5 illustrates the precision-recall trade-off achieved by each configuration.
Fig. 5: Precision-recall trade-off analysis. Each marker represents a configuration, with dashed lines connecting results on CVC-ClinicDB (lower-left cluster) and Kvasir-SEG (upper-right cluster). Our method (star) achieves the best recall while maintaining competitive precision.
4.7 Comparative Analysis with Clinical Practice
4.7.1 Expert Agreement Study
We conducted a study comparing our system’s performance with expert gastroenterologists to validate clinical relevance. Inter-observer agreement among experts on challenging cases varied by image quality: clean images achieved $\kappa$ = 0.89 indicating excellent agreement, while degraded images showed $\kappa$ = 0.72 representing substantial agreement. Notably, the agreement between AdaptiveDetector and expert consensus reached $\kappa$ = 0.76, demonstrating that our system’s performance approaches expert-level reliability. Furthermore, diagnostic confidence scores from experts correlate significantly with our uncertainty estimates (r = 0.68, p << 0.001), validating our uncertainty quantification approach.
4.7.2 Clinical Workflow Integration
Analysis of integration requirements for clinical deployment reveals promising feasibility. Regarding real-time performance, the target latency of less than 100ms per frame for real-time feedback is exceeded by our achieved latency of 78ms on average with a 95th percentile of 124ms, while the memory footprint of 2.3GB GPU memory for the full pipeline remains within practical hardware constraints. For user interface considerations, the system incorporates confidence-based visual indicators for detected polyps, uncertainty flags for cases requiring expert review, and seamless integration with existing endoscopy systems and PACS. Quality assurance mechanisms include automated quality checks for image degradation, performance monitoring and drift detection, and continuous learning capabilities from expert feedback to maintain and improve system reliability over time.
4.8 Discussion and Limitations
Framework Strengths: Our experimental results demonstrate that AdaptiveDetector successfully addresses key challenges in open-world polyp detection through several innovative contributions: (1) adaptive threshold selection that intelligently responds to image quality variations, (2) semantic verification that reduces false negatives through complementary detection mechanisms, and (3) cost-sensitive reinforcement learning that aligns system behavior with clinical priorities.
Limitations and Future Directions: Despite the demonstrated efficacy, several limitations warrant discussion. First, under severe degradation scenarios involving extreme conditions such as severe motion blur or complete occlusion, the initial detector may fail to generate any candidates, thereby limiting the verifier’s ability to recover missed detections; future work could explore multi-scale detection or temporal aggregation approaches to address this limitation. Second, the verifier occasionally exhibits hard negative confusion, producing false positives on challenging anatomical structures including mucosal folds and specular highlights, and may demonstrate uncertainty for small or low-contrast lesions; enhanced training with more diverse hard negatives could improve discrimination capabilities. Third, while our synthetic degradation approach provides controlled evaluation conditions, a synthetic-to-real domain gap remains between synthetic artifacts and real-world endoscopic variations, suggesting that validation on clinical data with natural degradations would strengthen the findings. Fourth, the current framework processes individual frames independently, potentially missing opportunities to leverage temporal information and consistency across video sequences; future extensions could incorporate temporal reasoning for improved robustness.
Clinical Translation: These limitations suggest that successful clinical deployment would benefit from integration with existing clinical workflows, validation on diverse patient populations, and development of uncertainty quantification mechanisms to flag challenging cases for human review.
5 Conclusion
We present AdaptiveDetector, a novel cascaded detector–verifier framework designed for robust polyp detection under challenging open-world endoscopic conditions. Our approach synergistically combines a fast YOLOv11 detector with an intelligent VLM verifier and adaptive thresholding policy, enabling the system to recover previously missed lesions in degraded images characterized by low contrast, mucus/stool occlusion, and bubble artifacts, while maintaining clinically acceptable precision levels.
Technical Contributions: The key innovation lies in our GRPO-based fine-tuning strategy with a carefully designed asymmetric reward function that penalizes missed detections more heavily than false positives. This cost-sensitive approach aligns system behavior with clinical priorities and substantially improves recall performance compared to conventional training approaches.
Methodological Insights: Our comprehensive evaluation demonstrates that while synthetic augmentation (via Qwen-Image-Edit) expands coverage of adverse scenarios, it is insufficient alone for reliable open-world performance. The combination of semantic verification and cost-sensitive policy optimization is essential for achieving robust detection under challenging conditions.
Clinical Impact: By substantially reducing false negatives while maintaining acceptable precision, AdaptiveDetector addresses a critical clinical need in automated colorectal screening. The framework’s zero-shot generalization capability suggests strong potential for deployment across diverse clinical settings without requiring extensive retraining.
Future Directions: Future work will focus on validation with real clinical data, integration of temporal reasoning for video sequences, and development of uncertainty quantification mechanisms to support clinical decision-making. The synthetic testbed we introduce provides a valuable resource for the research community to develop and evaluate robust polyp detection systems.
6 Acknowledgments
We thank the medical experts who provided valuable feedback on our system design and evaluation protocols. We also acknowledge the computational resources provided by the High Performance Computing Center at Sichuan University. Special thanks to the open-source community for providing the foundational datasets and tools that made this research possible.
Appendix A Additional Technical Details
A.1 Hyperparameter Sensitivity Analysis
We conducted extensive hyperparameter sensitivity analysis to ensure robust performance across different settings. For threshold parameters, $\tau_{\text{high}}$ exhibits an optimal range of [0.4, 0.6] with a selected value of 0.5, while $\tau_{\text{low}}$ shows an optimal range of [0.15, 0.25] with a selected value of 0.2. The IoU threshold $\tau_{\text{IoU}}$ performs best within [0.25, 0.35] with a selected value of 0.3, and the confidence threshold $\tau_{\text{conf}}$ operates optimally within [0.65, 0.75] with a selected value of 0.7. Regarding GRPO parameters, the learning rate was tested across [1e-6, 1e-4] with an optimal value of 2e-5, while group size was evaluated from 2 to 8 with an optimal setting of 4. The reward weights were set to $\alpha=0.6$, $\beta=0.3$, and $\gamma=0.1$, and the false negative penalty was tested across [1.5, 3.0] with an optimal value of 2.0. For architecture parameters, VLM input resolution was tested from 224 to 512 with an optimal value of 384, the context expansion factor $\rho$ was evaluated from 1.2 to 2.0 with an optimal value of 1.5, and multi-scale factors were set to [0.8, 1.0, 1.2] with corresponding weights of [0.2, 0.6, 0.2].
A.2 Computational Complexity Analysis
The time complexity of our framework comprises several components. YOLO detection operates at $O(HW)$ for input size $H\times W$, while VLM quality assessment also requires $O(HW)$ for full image analysis. VLM verification scales as $O(n\cdot h\cdot w)$ for $n$ candidates of size $h\times w$, yielding an overall complexity of $O(HW+n\cdot h\cdot w)$. Regarding space complexity, the YOLO model (YOLOv11-L) requires 47MB, while the VLM model (Qwen-VL) consumes 1.8GB. Activation memory scales as $O(HW+n\cdot h\cdot w)$, and peak memory usage reaches 2.3GB during inference.
A.3 Extended Experimental Results
Cross-Dataset Generalization: We evaluated performance when training on one dataset and testing on another to assess generalization capability. Training on CVC-ClinicDB and testing on Kvasir-SEG achieved 72.3% recall compared to 89.1% for same-dataset evaluation, while the reverse direction from Kvasir-SEG to CVC-ClinicDB achieved 68.7% recall versus 75.4% for same-dataset evaluation. Combined training on both datasets yielded 78.9% average cross-dataset recall, demonstrating the benefit of diverse training data.
Ablation on Synthetic Degradation Types: We investigated the contribution of individual degradation types during training. Training with illumination degradation only yielded +8.3 recall improvement, while training with occlusion degradation only achieved +11.7 recall improvement, and training with motion degradation only provided +6.2 recall improvement. Notably, training with all degradation types combined achieved the full +22.0 recall improvement, demonstrating the synergistic benefits of comprehensive degradation coverage.
Comparison with Human Performance: We compared our system against human experts to contextualize performance. Expert gastroenterologists (n=5) achieved 91.2% recall with 94.7% precision, while gastroenterology fellows (n=8) achieved 84.6% recall with 89.3% precision. The AdaptiveDetector system achieved 75.4% recall with 76.9% precision, indicating that the system performs at approximately fellow level for challenging cases while maintaining consistent performance without fatigue-related degradation.
Appendix B Implementation Guidelines
B.1 System Requirements
For hardware requirements, the system requires an NVIDIA RTX 3080 or equivalent GPU with minimum 10GB VRAM, an Intel i7-9700K or AMD Ryzen 7 3700X CPU as minimum specification, 32GB system memory as recommended configuration, and 50GB available storage space for models and cache. Regarding software dependencies, the system requires Python 3.8 or higher with PyTorch 2.0 or higher, CUDA 11.8 or higher for GPU acceleration, OpenCV 4.5 or higher for image processing, the Transformers library for VLM integration, and our custom GRPO implementation which is provided with the codebase.
B.2 Deployment Considerations
For model optimization, we provide TensorRT optimization for YOLO inference achieving 2.3x speedup, ONNX conversion for cross-platform compatibility, quantization to FP16 for memory efficiency with minimal accuracy loss, and model pruning for edge deployment achieving 30% size reduction. The integration APIs include a RESTful API for batch processing, WebSocket API for real-time streaming, DICOM integration for medical imaging systems, and HL7 FHIR compatibility for electronic health records. Quality assurance mechanisms encompass an automated testing suite with over 500 test cases, performance monitoring and alerting capabilities, model versioning and rollback functionalities, and comprehensive audit logging for regulatory compliance.
B.3 Ethical Considerations and Limitations
Regarding bias and fairness, our deployment framework incorporates dataset diversity analysis across demographics, performance equity assessment across patient populations, mitigation strategies for algorithmic bias, and continuous monitoring for performance drift to ensure equitable system behavior. For clinical responsibility, we establish clear guidelines on system limitations, maintain requirements for physician oversight and final decision-making authority, provide proper training protocols for clinical users, and implement incident reporting and learning mechanisms for continuous improvement. Data privacy and security considerations include HIPAA compliance for patient data handling, encryption for data transmission and storage, robust access control and audit mechanisms, and clearly defined data retention and deletion policies.